OpenStudy (anonymous):

can you simplify (x+1)/x?

OpenStudy (lgbasallote):

well...there is a way...

OpenStudy (lgbasallote):

\[\large \frac{(x+1)}{x} \implies \frac xx + \frac 1x \implies 1 + \frac 1x\] does that help?